Indexes & Trading

1. The price of Facebook stock is currently at $56.51
and you decide to buy 160 shares on margin. You borrow $1,500 from your broker and finance the remainder of the purchase with your own cash.

a. What is your margin (as a decimal value)?

b. If the price rises to $60, and the interest you
have to pay on the broker's loan is 3%, what is the net return (as a decimal value)?

c. If the broker's maintenance margin is 40%, what
is the minimum value that Facebook stock price can take before you are issued a margin call?

Ipdexes and Trading Indexes and Trading 1. The price of Facebook stock is currently at $56.51 and you decide to buy 160 shares on margin. You borrow $1,500 from your broker and finance the remainder of the purchase with your own cash. a. What is your margin (as a decimal value)? b. If the price rises to $60, and the interest you have to pay on the brokers loan is 3%, what is the net return (as a decimal value)? C. If the brokers maintenance margin is 40%, what is the minimum value that Facebook stock price can take before you are issued a margin call?
Answer #1

a) Margin = 160*56.5 - 1500 =  7540

b) Net Return = (60*160 - 1500*3% - 56.51*160)/7540 = 513.40/7540 = 6.81%

c) Margin Maintenance = 40% = ( 160*P - 1500)/160*P
64P = 160P - 1500
96P = 1500
Minimum Value or Price = 1500/96 = 15.63

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well